# Functional classification of metabolic networks

### v1

@@ Abstract
Complex chemical reaction networks are central to the function of living and non-living systems across a wide range of length scales, from microscopic organisms [1, 2, 3] and tissues [4, 5] to ecosystems [6, 7, 8] and planetary atmospheres [9, 10, 11]. Recent advances in experimental and computational methods have enabled the comprehensive reconstruction of metabolic processes in various biological systems [12, 13, 14, 15]. Similarly, the James Webb Space Telescope has produced high-quality spectroscopic data allowing the chemical characterization of exoplanet atmospheres in remarkable detail [16]. In bacterial communities, spatiotemporal pyruvate cross-feeding by swarming Bacillus subtilis has been observed; bacteria in the swarm front consume their preferred carbon source and deposit pyruvate which is consumed by bacteria in the bulk [17]. In mice and humans, models of metabolic processes have resolved metabolic cycles and energy use [5, 18, 19]. On the astrophysical scale, the topological distinctiveness of Earth’s atmosphere,from the atmospheres of other celestial bodies in the Solar System, has suggested the development of network-based biosignatures [11]. In these examples, the breadth of chemistries is shaped by processes that are potentially inaccessible to perturbation or measurement, such as evolution, cellular differentiation, and atmospheric development. The vast number of underlying degrees of freedom presents a challenge to the formation of methodical and functional comparisons between chemical reaction networks.
@@
In living systems, inferences of metabolic function from taxonomic classifications are inherently difficult; phylogenetically similar organisms may have vastly different metabolic capabilities [20, 21, 22, 23]. In complex organisms, tissues and organs share the same genetic code and are yet capable of various metabolic functions [4]. Hence, functional roles cannot be readily ascribed to organisms with similar genetic and evolutionary backgrounds. To tackle this problem, we introduce here a conceptual and computational framework for comparing the topologies of various chemical reaction networks, by measuring distances between the nullspaces of their stoichiometric matrices [24, 25, 26, 27, 28, 29].
Chemical reaction networks are naturally described by the mathematical theory of graphs in which chemical species are represented by vertices and physicochemical processes are represented by weighted and directed hyperedges which capture the direction and stoichiometric quantities of each process (Fig. 1) [26, 27]. Graphs of this flavor admit a matrix representation: the incidence or stoichiometric matrix $S$. Physically, these matrices satisfy the mass-action kinetic differential equation:
$\frac{\textnormal{d}c}{\textnormal{d}t}=Sv,$ (1)
where $c$ is a vector of concentrations and $v$ is a vector of fluxes. Each $v_{i}$ corresponds to the net rate of change in concentrations of the chemical species involved in process $i$. Each $v_{i}$ is a sum of directed fluxes: $v_{i}=v_{i}^{(+)}-v_{i}^{(-)}$ where directed fluxes are proportional to the probability of an encounter between reactants (or products) [28]. If the mathematical forms of these fluxes are known, the concentrations of chemicals in the network are readily obtained [30]. Without loss of generality, we will assume that all fluxes are reversible. Specifically, processes proceed in the forward direction if $v_{i}^{(+)}>v_{i}^{(-)}$, while for $v_{i}^{(+)}<v_{i}^{(-)}$ the process is reversed. The graph representation of chemical reaction networks makes distance metrics on graphs attractive choices. However, not all graph distance metrics are suitable for directed graphs with hyperedges; applications of these metrics typically ignore directionality or stoichiometry [31, 32]. Other existing metrics opt for computational tractability, such as reducing the scope to comparisons of the presence/absence of metabolic processes [33, 23] or feature vectors of topological measures derived from a graph-theoretic approach [34, 11]. We avoid these information losses by leveraging advances in parallel computing and numerical linear algebra to make calculations of stoichiometric nullspaces tractable for large datasets [35, 36]. At this point, one may ask why focus on nullspaces of stoichiometric matrices?
Principally, the stoichiometric nullspaces have natural physical interpretations in terms of the mass-action picture of Eq. 1. The right nullspace of a stoichiometric matrix satisfies the nonequilibrium steady-state flux condition required by flux balance‐based approaches [26, 27, 37], $Sv=0$. Namely, the right nullspace contains linear combinations of processes that result in net zero consumption and production of all chemical species. This is equivalent to currents satisfying Kirchoff’s current law in electrical circuits [28]. Contrast this with the left nullspace which consists of conservation laws for pools of chemicals [38, 28, 39, 40, 41]. Indeed, we see that for an element $w$ of the left nullspace:
@@
$d_{\text{Gr}(\infty,\infty)}(A,B)=\left(\left|k-\ell\right|\pi^{2}/4+\sum_{i=1}^{\min(k,\ell)}\theta_{i}^{2}\right)^{1/2},\ \ \ \Theta=\operatorname{diag}(\theta_{1},\ldots,\theta_{\min(k,\ell)})=\cos^{-1}\Sigma\ \ \ \text{where}\ \ \ A^{\intercal}B=U\Sigma V^{\intercal}.$ (2)
where the $\theta_{i}$ are the principal angles between the subspaces. If we consider $k\leq\ell$, the Grassmann distance metric provides a distance from the $\ell$-dimensional subspace $B$ to the furthest $\ell$-dimensional subspace that contains the $k$-dimensional subspace $A$. The symmetric statement is also true: it is the distance from the $k$-dimensional subspace $A$ to the furthest $k$-dimensional subspace contained in the $\ell$-dimensional subspace $B$ [29]. Herewithin, we will refer to the geodesic distance metric between right and left nullspaces, in the doubly infinite Grassmannian, as R-Grassmann and L-Grassmann, respectively.
We now move towards applications of the Grassmann distance to chemical reaction networks. First, we examine the effect of genetic perturbations in the genome of the model organism Escherichia coli K-12 MG1665 on these metabolic Grassmann distances [49]. We then consider genetically diverse organisms present in the AGORA2 (assembly of gut organisms through reconstruction and analysis, version 2) dataset which serves as a metabolic knowledge base for the human gut microbiome [49]. We show that the resulting Grassmann distances on nullspaces do not appreciably recapitulate taxonomic structures obtained from comparisons of the genetic background. To establish functional classifications, we identify metabolic process modalities that cluster organisms in each Grassmann distance, as well as the computationally tractable Jaccard distance [33, 23]. The final application of the Grassmann distance on the chemical reaction networks of human tissues [5] and planetary atmospheres [11] highlights the applicability of Grassmann distances to systems of different length scales and fields.
@@ Results
To illustrate the robustness of the Grassmann distances to genetic perturbations in the form of gene knockouts (KO), we use the AGORA2 metabolic network of the model organism E. coli K-12 MG1655. We perform single gene deletions on the genome-scale network by evaluating gene-protein-reaction (GPR) rules that describe the mapping from gene presence to reaction presence [49]. This allows us to obtain 389 unique networks corresponding to these deletions where a specific network may correspond to many single gene KOs. Of these KOs, $\sim$ 82% are computationally viable; they admit nonzero biomass flux in complete media with all possible external chemical inputs (SI). Furthermore, two-dimensional embeddings of the metabolic distance matrices, via stochastic gradient descent multidimensional scaling (sgd-MDS), demonstrate that the metrics considered here are capable of distinguishing computationally viable and unviable networks with the L-Grassmann distance attaining the best performance (SI).
@@
Metabolic processes are the smallest functional subunits of chemical reaction networks. As such, we seek to identify metabolic process modalities that lead to proximal nullspaces and opt for a cluster-based analysis, forgoing associating axes of low-dimensional embeddings to any particular aspect of metabolism. We use hierarchical clustering and tree-cuts with Ward’s linkage to analyze eight clusters for each metabolic distance. In principle, any desired level of granularity can be considered using this approach. In the spirit of phylogenetic taxonomy [57, 58], we examine the coarser cases of 2 and 4 clusters (SI).
@@
First, we note that $\sim$53of the networks considered here are either E. coli KOs or strains which contribute to the presence of primarily E. coli clusters. Both Grassmann distances produce two clusters with the same process attributed to them: R-Grassmann groups 1 & 7 and L-Grassmann groups 1 & (weakly) 2. By comparison to Fig. 2, we find that both groups 1, which are enriched in trans-cinnamate exchange, correspond to almost all of the E. coli K-12 MG1655 KO, whereas the remaining groups, which are depleted in ATP:dUMP phosphotransferase, are primarily composed of E. coli and Shigella flexneri strains. Let us now consider clusters on a metric-specific basis.
@@
R-Grassmann groups 5 and 6 correspond to processes with glucuronidated chemicals. Glucuronide moieties are added to metabolic substrates to increase hydrophilicity to faciliate elimination from the human body. As a result, glucuronidation is critical for the removal of unwanted endogenous molecules, drugs, and xenobiotics [50, 51, 52]. Group 8, which is enriched in the pyrophosphate:oxaloacetate carboxy-lyase reaction, is entirely composed of E. coli strains. This reaction produces inorganic phosphate, carbon dioxide, and phosphoenolpyruvate from pyrophosphate (PPi) and oxaloacetate. This process is analogous to the gluconeogeneic PECK reaction, which uses ATP instead of PPi as a phosphate donor, but is unlikely to share an evolutionary origin [60, 61, 62].
@@
Figure 5: Hierarchical clustering of metabolic distances reveals functional groups of organisms enriched or depleted in specific metabolic processes. Multidimensional scaling embeddings of the metabolic distance matrices are shown colored by clusters obtained by hierarchical clustering of the matrices with Ward linkage and a tree-cut to produce 8 disjoint clusters. We use normalized variation of information (nVI) to assess the validity of assignment to metabolic processes: zero nVI corresponds to perfect matching, whereas unity nVI corresponds to maximally distinct matching. Mean silhouette scores $\langle s\rangle$ are computed for each distance using the eight disjoint clusters as cluster assignments. Abbreviations: ABC = ATP-binding cassette transporters, ATP = adenosine triphosphate, CDP = cytosine diphosphate, CTP = cytosine triphosphate, F6P = fructose-6-phosphate, ITP = inosine triphosphate, SAH = S-adenosylhomocysteine, T6P = tagatose-6-phosphate, dUMP = deoxyuridine monophosphate.
@@
Figure 6: Comparisons of human tissues and planetary atmosphere highlights applicability of metabolic Grassmann distances to other complex chemical reaction networks. (A) Multidimensional scaling embeddings of metabolic Grassmann distance computed on sex-specific human tissue and organ metabolic networks are shown colored by 8 clusters obtained by hierarchical clustering of the matrices with Ward linkage and a tree-cut to produce 8 disjoint clusters. The most distinct organs in the right nullspace are the kidneys, liver, colon, and brain, whereas for the left nullspace, we find that the pancreas, colon, lungs, and brain are distal. Graphical elements were adapted from Servier Medical Art under a Creative Commons Attribution 4.0 license. (B) Multidimensional scaling embeddings of Metabolic Grassmann distance computed on 8 published planetary atmosphere chemical reaction networks serve as a proof of concept based on limited data availability. As additional planetary networks becomes available, these can be incorporated. Distance matrix labels are in order of distance from the Sun. Human tissue and planetary atmosphere chemical reaction networks were obtained from [5] and [11], respectively. Planetary graphical elements were derived from photographs obtained from the National Aeronautics and Space Administration.
@@
Despite metabolic differences between tissues of different sexes, we find that the Grassmann distances group networks by tissue type (Fig. 6A). Moreover, we observe clustering based on hematopoietic-stem cell origin (B cells, CD4 T cells, monocytes, NK cells, platelets, and red blood cells) [69]. We observe that for the right nullspace, the kidneys, liver, colon, and brain remain distinguishable from most networks in a sgd-MDS embedding of the distance matrix, whereas for the left nullspace, we find that the pancreas, colon, lungs, and brain are the distal tissues (Fig. 6A). This highlights the tissue-specific diversity of steady-state fluxes and conservation laws.
To move towards chemical reaction networks on the astrophysical scale, we consider eight atmospheric networks corresponding to six celestial bodies (Fig. 6B). The networks used here are derived from a study on the chemical reaction networks of planetary atmospheres using graph topological measures [11]. With the exception of the “textbook” Earth model [10], these networks were obtained from published studies simulating the photochemistry of atmospheres using the photochemistry–transport model KINETICS from Caltech–JPL [9]. Importantly, we observe that the most distal network corresponds to the “textbook” Earth atmospheric model, which suggests network curatorial effects on Grassmann distances. We are aware that data on planetary chemical reaction networks are currently limited, but as more become available these can be incorporated.
@@ Discussion
For the microbial networks considered here, the processes identified for R-Grassmann group 5 and Jaccard groups 1 & 7 occur across the periplasm, the region between the outer and inner cell membrane in bacteria with two membranes (diderms) or the region between the cell membrane and the cell well in bacteria with a single membrane (monoderms) [77, 78, 79, 80, 81]. In general, monoderms are Gram-positive and diderms are Gram-negative; however, there are notable exceptions [80, 81]. Previous work has shown that Gram stain reactivity of organisms produces well-defined clusters with the Jaccard distance [23]. We recapitulate this result with the Jaccard distance which outperforms the Grassmann distances in this regard (SI). We note that the Gram-negative networks in mini-AGORA2 have larger numbers of metabolic processes and metabolites than their Gram-positive counterparts (SI). Is it then surprising that the Jaccard distance, a metric based on set overlap and size, produces well-defined clusters on the basis of Gram-stain reactivity? We also note that AGORA2 networks were curated to include drug biotransformation and degradation reactions to enable the modeling of personalized gut microbial drug metabolism [49]. That we appreciably observe this feature when clustering with the R-Grassmann and Jaccard distances further underscores the need for a nuanced and context-aware interpretation of metabolic distances.
@@
Until now, we have considered open chemical reaction networks: systems forced away from thermodynamic equilibrium by the exchange of mass and energy with the environment. These exchanges are described by fluxes across the system boundary and correspond to columns of the stoichiometric matrix with all positive or negative entries (Fig. 1C). In open chemical reaction networks, the number of conservation laws observed cannot be larger than the number observed by their closed counterparts. As such, while closed chemical reaction networks realize nontrivial left nullspaces, the does not always apply to about open networks [38, 28] (SI). Considering the form of the Grassmann distance metric (Eq. 2), L-Grassmann distances on closed networks will necessarily impose an upper bound on their open network counterparts. We notably observe this in the comparison of the planetary atmosphere networks, where four networks have a trivial left nullspace and consequently map to the same point in a sgd-MDS embedding (Fig. 6B). Consistent with an increase in the number of conservation laws, this degeneracy is remedied by the removal of exchange processes (SI). Although we also observe changes to the L-Grassmann distances of the microbial and human tissue networks (SI), we note that biological systems are phenomonologically open. Energetic and material demands for maintenance and growth are satisfied by the efflux of chemicals from the environment [28]. We do not address here the extent to which a chemical reaction network is open (or closed) at any given point in time. Instead, we take a conservative approach in using open networks as removing any exchange process with the environment may introduce additional conservation laws which are not guaranteed to hold for all time.

### v2

@@ Abstract
Complex chemical reaction networks are central to the function of living and non-living systems across a wide range of length scales, from microscopic organisms [1, 2, 3] and tissues [4, 5] to ecosystems [6, 7, 8, 9] and planetary atmospheres [10, 11, 12]. Recent advances in experimental and computational methods have enabled the comprehensive reconstruction of metabolic processes in various biological systems [13, 14, 15, 16]. In bacterial communities, spatiotemporal pyruvate cross-feeding by swarming Bacillus subtilis has been observed; bacteria in the swarm front consume their preferred carbon source and deposit pyruvate which is consumed by bacteria in the bulk [17]. In mice and humans, models of metabolic processes have resolved metabolic cycles and energy use [5, 18, 19]. On the astrophysical scale, the distinctiveness of Earth’s atmosphere, from the atmospheres of other celestial bodies in the Solar System, has suggested the development of network-based biosignatures [12]. The James Webb Space Telescope and other sources have produced high-quality spectroscopic data that will allow for the chemical characterization of exoplanet atmospheres in remarkable detail [20, 21]. Across all these examples, the breadth of chemistries is shaped by processes that are potentially inaccessible to perturbation or measurement, such as evolution, cellular differentiation, and atmospheric development. The vast number of underlying degrees of freedom presents a challenge to the formation of methodical and functional comparisons between chemical reaction networks.
In living systems, inferences of metabolic function from taxonomic classifications are inherently difficult; phylogenetically similar organisms may have vastly different metabolic capabilities [22, 23, 24, 25]. In complex organisms, tissues and organs share the same genetic code and are yet capable of various metabolic functions [4]. Hence, functional roles cannot be readily ascribed to organisms with similar genetic and evolutionary backgrounds. To tackle this problem, we introduce here a conceptual and computational framework for comparing the topologies of various chemical reaction networks, by measuring distances between the nullspaces of their stoichiometric matrices [26, 27, 28, 29, 30, 31], which encode the steady-state network fluxes and fundamental conservation laws.
Chemical reaction networks are naturally described by graphs in which chemical species are represented by vertices and physicochemical processes are represented by weighted and directed hyperedges which capture the direction and stoichiometric quantities of each reaction process (Fig. 1) [28, 29]. Graphs of this flavor admit a matrix representation: the incidence or stoichiometric matrix $S$. Physically, these matrices satisfy the mass-action kinetic differential equation:
$\frac{\textnormal{d}c}{\textnormal{d}t}=Sv,$ (1)
where $c$ is a vector of concentrations and $v$ is a vector of fluxes. Each $v_{i}$ corresponds to the net rate of change in concentrations of the chemical species involved in process $i$. Each $v_{i}$ is a sum of directed fluxes: $v_{i}=v_{i}^{(+)}-v_{i}^{(-)}$ where directed fluxes are proportional to the probability of an encounter between reactants (or products) [30]. If the mathematical forms of these fluxes are known, the concentrations of chemicals in the network are readily obtained [32]. Without loss of generality, we will assume that all fluxes are reversible. Specifically, processes proceed in the forward direction if $v_{i}^{(+)}>v_{i}^{(-)}$, while for $v_{i}^{(+)}<v_{i}^{(-)}$ the process is reversed. We note that many complex nonlinear systems can be recast as linear systems of the form of Eq. (1) where the nonlinear details are absconded in flux-like functions [33].
@@
The graph representation of chemical reaction networks makes distance metrics on graphs attractive choices. However, not all graph distance metrics are suitable for directed graphs with hyperedges; applications of these metrics typically ignore directionality or stoichiometry [34, 35]. Other existing metrics opt for computational tractability, such as reducing the scope to comparisons of the presence/absence of metabolic processes [36, 25] or feature vectors of topological measures derived from a graph-theoretic approach [37, 12]. We avoid these information losses by leveraging advances in parallel computing and numerical linear algebra to make calculations of stoichiometric nullspaces tractable for large datasets [38, 39]. At this point, one may ask why focus on nullspaces of stoichiometric matrices?
Principally, the stoichiometric nullspaces have clear physical interpretations in terms of the mass-action picture of Eq. 1. The right nullspace of a stoichiometric matrix satisfies the nonequilibrium steady-state flux condition required by flux balance‐based approaches [28, 29, 40], $Sv=0$. Namely, the right nullspace contains linear combinations of processes that result in net zero consumption and production of all chemical species. This is equivalent to currents satisfying Kirchoff’s current law in electrical circuits [30]. Contrast this with the left nullspace which consists of conservation laws for pools of chemicals [41, 30, 42, 43, 44]. Indeed, we see that for an element $w$ of the left nullspace:
@@
$d_{\text{Gr}(\infty,\infty)}(A,B)=\left(\left|k-\ell\right|\pi^{2}/4+\sum_{i=1}^{\min(k,\ell)}\theta_{i}^{2}\right)^{1/2}$ (2)
If we consider $k\leq\ell$, the Grassmann distance metric provides a distance from the $\ell$-dimensional subspace $B$ to the furthest $\ell$-dimensional subspace that contains the $k$-dimensional subspace $A$. The symmetric statement is also true: it is the distance from the $k$-dimensional subspace $A$ to the furthest $k$-dimensional subspace contained in the $\ell$-dimensional subspace $B$ [31]. Herewithin, we will refer to the geodesic distance metric between right and left nullspaces, in the doubly infinite Grassmannian, as R-Grassmann and L-Grassmann, respectively.
We now move towards applications of the Grassmann distance to chemical reaction networks. First, we examine the effect of genetic perturbations in the genome of the model organism Escherichia coli K-12 MG1665 on these metabolic Grassmann distances [52]. We then consider genetically diverse organisms present in the AGORA2 (assembly of gut organisms through reconstruction and analysis, version 2) dataset which serves as a metabolic knowledge base for the human gut microbiome [52]. We show that the resulting Grassmann distances on nullspaces can differ substantially from taxonomic structures obtained from comparisons of the genetic background. To establish functional classifications, we identify metabolic process modalities that cluster organisms in each Grassmann distance, as well as the computationally tractable Jaccard distance [36, 25]. The final application of the Grassmann distance on the chemical reaction networks of human tissues [5] and planetary atmospheres [12] highlights the applicability of Grassmann distances to systems of different length scales and fields.
@@ Results
To illustrate the robustness of the Grassmann distances to genetic perturbations in the form of gene knockouts (KO), we use the AGORA2 metabolic network of the model organism E. coli K-12 MG1655. We perform single gene deletions on the genome-scale network by evaluating gene-protein-reaction (GPR) rules that describe the mapping from gene presence to reaction presence [52]. This allows us to obtain 389 unique networks corresponding to these deletions where a specific network may correspond to many single gene KOs. Of these KOs, $\sim$ 82% are computationally viable; they admit nonzero biomass flux in complete media with all possible external chemical inputs (SI Appendix). Furthermore, two-dimensional embeddings of the metabolic distance matrices, via stochastic gradient descent multidimensional scaling (sgd-MDS), demonstrate that the metrics considered here are capable of distinguishing computationally viable and unviable networks with the L-Grassmann distance attaining the best performance (SI Appendix).
@@
Metabolic processes are the smallest functional subunits of chemical reaction networks. As such, we seek to identify metabolic process modalities that lead to proximal nullspaces and opt for a cluster-based analysis, forgoing associating axes of low-dimensional embeddings to any particular aspect of metabolism. We use hierarchical clustering and tree-cuts with Ward’s linkage to analyze eight clusters for each metabolic distance. In principle, any desired level of granularity can be considered using this approach. In the spirit of phylogenetic taxonomy [60, 61], we touch upon coarser cases of 2 and 4 clusters (SI Appendix).
@@
First, we note that $\sim$53of the networks considered here are either E. coli KOs or strains which contribute to the presence of primarily E. coli clusters. Both Grassmann distances produce two clusters with the same process attributed to them: R-Grassmann groups 1 & 7 and L-Grassmann groups 1 & (weakly) 2. By comparison to Fig. 2, we find that both groups 1 correspond to almost all of the E. coli K-12 MG1655 KO, whereas the remaining groups are primarily composed of E. coli and Shigella flexneri strains. Let us now consider clusters on a metric-specific basis.
@@
R-Grassmann groups 5 and 6 correspond to processes with glucuronidated chemicals. Glucuronide moieties are added to metabolic substrates to increase hydrophilicity to facilitate elimination from the human body. As a result, glucuronidation is critical for the removal of unwanted endogenous molecules, drugs, and xenobiotics [53, 54, 55]. Group 8, which is enriched in the pyrophosphate:oxaloacetate carboxy-lyase reaction, is entirely composed of E. coli strains. This reaction produces inorganic phosphate, carbon dioxide, and phosphoenolpyruvate from pyrophosphate (PPi) and oxaloacetate. This process is analogous to the gluconeogeneic PECK reaction, which uses ATP instead of PPi as a phosphate donor, but is unlikely to share an evolutionary origin [63, 64, 65].
@@
Figure 5: Hierarchical clustering of metabolic distances reveals functional groups of organisms enriched or depleted in specific metabolic processes. Multidimensional scaling embeddings of the metabolic distance matrices are shown colored by clusters obtained by hierarchical clustering of the matrices with Ward linkage and a tree-cut to produce 8 disjoint clusters. We use normalized variation of information (nVI) to assess the validity of assignment to metabolic processes: zero nVI corresponds to perfect matching, whereas unity nVI corresponds to maximally distinct matching. Mean silhouette scores $\langle s\rangle$ are computed for each distance using the eight disjoint clusters as cluster assignments. Abbrevations: ABC = ATP-binding cassette transporters, ATP = adenosine monophosphate, CDP = cytosine diphosphate, CoA = coenzyme A, CTP = cytosine triphosphate, F6P = fructuse-6-phosphate, ITP = inosine triphosphate, SAH = S-adenosylhomocysteine, T6P = tagatose-6-phosphate, dUMP = deoxyuridine monophosphate.
@@
Figure 6: Comparisons of human tissues and planetary atmosphere highlights applicability of metabolic Grassmann distances to other complex chemical reaction networks. (A) Multidimensional scaling embeddings of metabolic Grassmann distance computed on sex-specific human tissue and organ metabolic networks are shown colored by 8 clusters obtained by hierarchical clustering of the matrices with Ward linkage and a tree-cut to produce 8 disjoint clusters. The most distinct organs in the right nullspace are the kidneys, liver, colon, and brain, whereas for the left nullspace, we find that the pancreas, colon, lungs, and brain are distal. Graphical elements were adapted from Servier Medical Art under a Creative Commons Attribution 4.0 license. (B) Multidimensional scaling embeddings of Metabolic Grassmann distance computed on 8 published planetary atmosphere chemical reaction networks serve as a proof of concept based on limited data availability. As additional planetary networks becomes available, these can be incorporated. Distance matrix labels are in order of distance from the Sun. Human tissue and planetary atmosphere chemical reaction networks were obtained from [5] and [12], respectively. Planetary graphical elements were obtained from the National Aeronautics and Space Administration.
@@
Despite metabolic differences between tissues of different sexes, we find that the Grassmann distances group networks by tissue type (Fig. 6A). We also observe clustering based on hematopoietic-stem cell origin (B cells, CD4 T cells, monocytes, NK cells, platelets, and red blood cells) [72]. We observe that for the right nullspace, the kidneys, liver, colon, and brain remain distinguishable from most networks in a sgd-MDS embedding of the distance matrix, whereas for the left nullspace, we find that the pancreas, colon, lungs, and brain are the distal tissues (Fig. 6A).
To move towards chemical reaction networks on the astrophysical scale, we consider eight atmospheric networks corresponding to six celestial bodies (Fig. 6B). The networks used here are derived from a study on the chemical reaction networks of planetary atmospheres using graph topological measures [12]. With the exception of the “textbook” Earth model [11], these networks were obtained from published studies simulating the photochemistry of atmospheres using the photochemistry–transport model KINETICS from Caltech–JPL [10]. Importantly, we observe that the most distal network corresponds to the “textbook” Earth atmospheric model, which suggests network curatorial effects on Grassmann distances. We note that data on planetary chemical reaction networks are currently limited, but as more become available these can be readily incorporated.
@@ Discussion
For the microbial networks considered here, the processes identified for R-Grassmann group 5 and Jaccard groups 1 & 7 occur across the periplasm, the region between the outer and inner cell membrane in bacteria with two membranes (diderms) or the region between the cell membrane and the cell well in bacteria with a single membrane (monoderms) [80, 81, 82, 83, 84]. In general, monoderms are Gram-positive and diderms are Gram-negative; however, there are notable exceptions [83, 84]. Previous work has shown that Gram stain reactivity of organisms produces well-defined clusters with the Jaccard distance [25]. We recapitulate this result with the Jaccard distance which outperforms the Grassmann distances in this regard (SI Appendix). We note that the Gram-negative networks in mini-AGORA2 have larger numbers of metabolic processes and metabolites than their Gram-positive counterparts (SI Appendix). Is it then surprising that the Jaccard distance, a metric based on set overlap and size, produces well-defined clusters on the basis of Gram-stain reactivity? We also note that AGORA2 networks were curated to include drug biotransformation and degradation reactions to enable the modeling of personalized gut microbial drug metabolism [52]. That we appreciably observe this feature when clustering with the R-Grassmann and Jaccard distances further underscores the need for a nuanced and context-aware interpretation of metabolic distances.
@@
Until now, we have considered open chemical reaction networks: systems forced away from thermodynamic equilibrium by the exchange of mass and energy with the environment. These exchanges are described by fluxes across the system boundary and correspond to columns of the stoichiometric matrix with all positive or negative entries (Fig. 1C). In open chemical reaction networks, the number of conservation laws observed cannot be larger than the number observed by their closed counterparts. As such, while closed chemical reaction networks realize nontrivial left nullspaces, the same does not always apply to open networks [41, 30] (SI Appendix). We find that L-Grassmann distances on closed networks generally exceed their open network counterparts (SI Appendix). We notably observe this in the comparison of the planetary atmosphere networks, where four networks have a trivial left nullspace and consequently map to the same point in a sgd-MDS embedding (Fig. 6B). Consistent with an increase in the number of conservation laws, this degeneracy is remedied by the removal of exchange processes (SI Appendix). Although we also observe changes to the L-Grassmann distances of the microbial and human tissue networks (SI Appendix), we note that biological systems are phenomonologically open. Energetic and material demands for maintenance and growth are satisfied by the efflux of chemicals from the environment [30]. We do not address here the extent to which a chemical reaction network is open (or closed) at any given point in time. Instead, we take a conservative approach in using open networks as removing any exchange process with the environment may introduce additional conservation laws which are not guaranteed to hold for all time.